# From Text to Multimodality: Exploring the Evolution and Impact of Large Language Models in Medical Practice

### v5

@@ V Future Directions and Conclusion
This review has emphasized the potential of LLMs and MLLMs to revolutionize medicine and healthcare. While showing early promise in areas like patient-trial matching [108], generating radiology reports [109], and assisting with clinical diagnostics [62], LLMs are still in their early stages. Significant research gaps remain and must be addressed to unlock their full potential and ensure safe, responsible, and equitable integration into clinical practice.
@@ V-A Data Augmentation and Access
Data Augmentation and Access: The scarcity of large-scale, high-quality, and diverse multimodal datasets is a major bottleneck [106]. This is especially true for languages other than English [32]. Future research shall focus on:
Dataset Creation and Curation: Developing large, well-annotated datasets encompassing diverse medical specialties, patient populations, and languages is crucial [85]. This includes incorporating visual data like medical images, alongside text from EHRs, clinical notes, and medical literature [110]. Datasets should represent real-world scenarios and address issues like imbalanced data [38]. The OmniMedVQA benchmark provides a good example of a comprehensive dataset that addresses some of these challenges [48].
Privacy-Preserving Data Sharing: Investigating innovative methods like federated learning [33] to enable collaborative data sharing and model training while preserving patient privacy.
Standardization and Interoperability: Developing standardized data formats and categories to facilitate data integration and interoperability across different healthcare systems and institutions. This is crucial for training models that can generalize well to new settings [111].
@@ V-B Advanced Modality Alignment
A key area for future research is developing more sophisticated methods for aligning different modalities. Future research could focus on developing novel architectures and training strategies that can better capture the complex relationships between different modalities, leading to more accurate and robust predictions [112].
@@ V-C Interpretability and Explainability
A critical area for future research is enhancing the interpretability and explainability of MLLMs. This lack of transparency in current LLMs MLLMs can hinder trust and adoption in clinical settings. Traditional evaluation methods for MLLMs are usually insufficient for clinical settings, as they don’t adequately assess their impact on real-world workflows [73]. Future research should focus on developing methods to make MLLM decision-making processes more transparent and understandable, such as generating human-readable explanations for their predictions or visualizing the processes that contribute to their decisions.
@@ V-D Robust Evaluation Frameworks
Robust and standardized evaluation frameworks becomes increasingly critical as MLLMs become increasingly sophisticated. Current evaluation methods often rely on limited datasets and metrics to non-clinical tasks, restricting the potential to capture the full range of capabilities and biases [113]. To ensure the safe and effective of clinical MLLMs, future effort should spend on developing additional standardized benchmarks with closer clinical relevance.
@@ V-E Ethics and Compliance
The ethical implications of MLLMs in healthcare cannot be overstated. Its training data contain sensitive patient information, raising concerns about privacy and data security [29]. Biased training data can lead to discriminatory outcomes, potentially exacerbating existing health disparities. Therefore, clear regulatory frameworks and guidelines are necessary to govern the development, deployment, and use of MLLMs in clinical settings [35]. Addressing these ethical and compliance challenges will be beneficial to establish trust and ensure the responsible use of MLLMs in healthcare.

### v6

@@ V Future Directions and Conclusion
This review has emphasized the potential of LLMs and MLLMs to revolutionize medicine and healthcare. While showing early promise in areas like patient-trial matching [Dyke2024Autonomous], generating radiology reports [Liu2023Radiology], and assisting with clinical diagnostics [Tu2023Towards], LLMs are still in their early stages. Significant research gaps remain and must be addressed to unlock their full potential and ensure safe, responsible, and equitable integration into clinical practice.
@@ V-A Data Augmentation and Access
Data Augmentation and Access: The scarcity of large-scale, high-quality, and diverse multimodal datasets is a major bottleneck [Chunyuan2023LLaVA]. This is especially true for languages other than English [Iker2024MedMT5]. Future research shall focus on:
Dataset Creation and Curation: Developing large, well-annotated datasets encompassing diverse medical specialties, patient populations, and languages is crucial [Qiu2024Towards]. This includes incorporating visual data like medical images, alongside text from EHRs, clinical notes, and medical literature [Xie2024Me]. Datasets should represent real-world scenarios and address issues like imbalanced data [Meng2024application]. The OmniMedVQA benchmark provides a good example of a comprehensive dataset that addresses some of these challenges [Yutao2024OmniMedVQA].
Privacy-Preserving Data Sharing: Investigating innovative methods like federated learning [Wang2023Accelerating] to enable collaborative data sharing and model training while preserving patient privacy.
Standardization and Interoperability: Developing standardized data formats and categories to facilitate data integration and interoperability across different healthcare systems and institutions. This is crucial for training models that can generalize well to new settings [Longwell2024Performance].
@@ V-B Advanced Modality Alignment
A key area for future research is developing more sophisticated methods for aligning different modalities. Future research could focus on developing novel architectures and training strategies that can better capture the complex relationships between different modalities, leading to more accurate and robust predictions [Meiqi2024Quantifying].
@@ V-C Interpretability and Explainability
A critical area for future research is enhancing the interpretability and explainability of MLLMs. This lack of transparency in current LLMs MLLMs can hinder trust and adoption in clinical settings. Traditional evaluation methods for MLLMs are usually insufficient for clinical settings, as they don’t adequately assess their impact on real-world workflows [Mehandru2024Evaluating]. Future research should focus on developing methods to make MLLM decision-making processes more transparent and understandable, such as generating human-readable explanations for their predictions or visualizing the processes that contribute to their decisions.
@@ V-D Robust Evaluation Frameworks
Robust and standardized evaluation frameworks becomes increasingly critical as MLLMs become increasingly sophisticated. Current evaluation methods often rely on limited datasets and metrics to non-clinical tasks, restricting the potential to capture the full range of capabilities and biases [Huang2024Comprehensive]. To ensure the safe and effective of clinical MLLMs, future effort should spend on developing additional standardized benchmarks with closer clinical relevance.
@@ V-E Ethics and Compliance
The ethical implications of MLLMs in healthcare cannot be overstated. Its training data contain sensitive patient information, raising concerns about privacy and data security [peng2024securinglargelanguagemodels]. Biased training data can lead to discriminatory outcomes, potentially exacerbating existing health disparities. Therefore, clear regulatory frameworks and guidelines are necessary to govern the development, deployment, and use of MLLMs in clinical settings [Ong2024Ethical]. Addressing these ethical and compliance challenges will be beneficial to establish trust and ensure the responsible use of MLLMs in healthcare.